Please help, I don't even know where to start on this one......Thank you!



Taylor has 15 bills of $5 dollars and $10 dollars. Their total value is $105 dollars. Find the number of the $5 dollar and $10 dollar bills that Taylor has.

Answer by mananth(16949)

Let the number of $ 5 bills be x
the number of $10 bills will be 15-x
5x+10(15-x)=105
5x+150-10x=105
-5x-45
x=-45/-5
x= 9 number of $ 5 bills
balance $10 bills are 6 in number